OpenStudy (anonymous):

What is the lateral area of a triangular prism with base edges all equaling 4 feet and a height of 10 feet? 160 ft2 180 ft2 140 ft2 120 ft2

OpenStudy (anonymous):

Area of each triangular face = 4*sqrt(3) sq ft = 13.86 sq ft (to 2 dp) Area of each rectangular face = 40 sq ft Total area = 2*13.86 + 3*40 = 147.71 sq ft

OpenStudy (nottim):

dont see what you need here

OpenStudy (experimentx):

120

OpenStudy (experimentx):

you should do some reading, you cannot always depend on other for answers.

OpenStudy (anonymous):

that exactly what i thought but i wasnt sure

Directrix (directrix):

Latereal Area = p h where p is the perimeter of the base LA = (4+4+4) *10 LA = 12*10 LA = 120 square feet